# A Tutorial on MIMO-OFDM ISAC: From Far-Field to Near-Field

### v1

@@ Abstract
Integrated sensing and communication (ISAC) is one of the key usage scenarios for future sixth-generation (6G) mobile communication networks, where communication and sensing (C&S) services are simultaneously provided through shared wireless spectrum, signal processing modules, hardware, and network infrastructure. Such an integration is strengthened by the technology trends in 6G, such as denser network nodes, larger antenna arrays, wider bandwidths, higher frequency bands, and more efficient utilization of spectrum and hardware resources, which incentivize and empower enhanced sensing capabilities. Moreover, emerging applications such as Internet-of-Everything (IoE), autonomous ground and aerial vehicles, virtual reality/augmented reality (VR/AR), and connected intelligence have intensified the demands for both high-quality C&S services, accelerating the development and implementation of ISAC in wireless networks. As the dominant waveform used in contemporary communication systems, orthogonal frequency division multiplexing (OFDM) is still expected to be a very competitive technology for 6G, rendering it necessary to thoroughly investigate the potential and challenges of OFDM ISAC. Thus, this paper aims to provide a comprehensive tutorial overview of ISAC systems enabled by large-scale multi-input multi-output (MIMO) and OFDM technologies and to discuss their fundamental principles, advantages, and enabling signal processing methods. To this end, a unified MIMO-OFDM ISAC system model is first introduced, followed by four frameworks for estimating parameters across the spatial, delay, and Doppler domains, including parallel one-domain, sequential one-domain, joint two-domain, and joint three-domain parameter estimation. Next, sensing algorithms and performance analyses are presented in detail for far-field scenarios where uniform plane wave (UPW) propagation is valid, followed by their extensions to near-field scenarios where uniform spherical wave (USW) characteristics need to be considered. Finally, this paper points out open challenges and outlines promising avenues for future research on MIMO-OFDM ISAC.
Index Terms:
Integrated sensing and communication (ISAC), MIMO-OFDM ISAC, far-field, near-field, super resolution.
@@ I Introduction
ISAC originated from radar-communication coexistence (RCC) studies [13], which initially focused on interference mitigation for collocated radar and communication systems. With the evolution of wireless communication and digital radar technologies, such as multi-input multi-output (MIMO) [14, 15] and orthogonal frequency division multiplexing (OFDM) [16], the gap between C&S has been gradually bridged. This technological convergence drove the emergence of dual functional radar communication (DFRC) systems [17, 18] that achieve higher spectral efficiency (SE) and hardware integration. To enable the coexistence of radar and communication systems within the same frequency band, a portion of the sub-6 GHz bands was released for radar and communication [19], which promoted the evolution of DFRC towards ISAC. Afterwards, extensive research efforts have been devoted to optimizing ISAC performance by virtue of information theory [20, 21, 22], waveform design [23, 24, 25, 26, 27, 28, 29], signal processing techniques [30, 31], and experiments and prototyping [32, 33].
@@
There are three typical design paradigms for ISAC, namely, communication-centric design, sensing-centric design, and co-design of C&S. In this paper, we focus on communication-centric designs that exploit communication waveforms and design frameworks to achieve sensing functions. Leveraging globally and densely deployed wireless infrastructure and networks, communication-centric ISAC has emerged as one of the most promising approaches for 6G systems, providing significant advantages over separate C&S designs as follows.
@@
$\bullet$ Efficient spectrum utilization: The massive number of Internet-of-Things (IoT) devices and soaring demands for high-speed data services have imposed an unprecedented strain on the limited spectrum resources available for C&S [34]. In conventional DFRC systems, C&S are performed in non-overlapping frequency bands, resulting in fragmented and poor utilization of spectrum resources. Communication-centric ISAC provides an efficient means to greatly improve spectrum efficiency, since the spectrum pool can be flexibly allocated and even completely reused by C&S tasks, as will be discussed in Section II-B.
@@
$\bullet$ Pervasive sensing via wireless communication networks: Pervasive sensing is an important ISAC vision for 6G that cannot be achieved by traditional radar sensing networks with limited sensing coverage. This issue can be resolved by leveraging existing wireless infrastructure, including cellular base stations (BSs) and backhaul links that are globally and densely deployed. Working cooperatively, a collection of wireless networks can form a powerful and pervasive sensor network with ubiquitous coverage and connectivity.
@@
$\bullet$ Enhanced cost/energy efficiency: Enhancing energy efficiency (EE) and cost efficiency (CE) is an important goal for 6G and ISAC. Pervasive sensing that relies on conventional standalone radars would require the deployment of a large amount of radar equipment, thus greatly degrading the EE/CE. Instead, leveraging existing communication systems to perform wireless sensing can significantly reduce the hardware cost and energy consumption for sensing, bringing great value to 6G.
@@
$\bullet$ High SE: OFDM systems employ inverse fast Fourier transform (IFFT)/ fast Fourier transform (FFT) operations to achieve frequency domain transformations, dividing the channel bandwidth into multiple orthogonal subcarriers for parallel data transmission, thus significantly improving SE.
@@
$\bullet$ Robustness against ISI: The use of a cyclic prefix (CP) in OFDM is a pivotal technique that can efficiently mitigate the impact of ISI caused by multipath.
@@
$\bullet$ Easy equalization: Since OFDM equalization is performed in the frequency domain, the receiver only needs to perform one-tap equalization on each subcarrier, making it much simpler and more efficient than time-domain equalization in single-carrier systems.
@@
$\bullet$ Flexible time-frequency resource allocation: The number and spacing of subcarriers, together with the power and bit loading to each subcarrier can be dynamically adjusted based on available channel state information (CSI), thereby maximizing network throughput and minimizing error rates while ensuring desired performance for individual users.
@@
For target sensing, OFDM has also been considered as a promising waveform with the following advantages.
@@
$\bullet$ Array-manifold-like structure in the subcarrier and symbol domains: After removing random information-bearing symbols, the remaining OFDM signal exhibits an array-manifold-like structure in both the subcarrier and symbol domains, enabling efficient estimation of the delay and Doppler frequency for the sensing targets, as will be shown in detail in Section IV-B.
@@
$\bullet$ Efficient decoupled estimation of delay and Doppler: The phase variations in OFDM channels induced by delay and Doppler in the subcarrier and symbol domains are independent. This allows for efficient decoupled estimation of delay and Doppler, where the data in the symbol domain can be used as snapshots for delay estimation and vice versa.
@@
$\bullet$ “Thumbtack-shaped” ambiguity function: In both the subcarrier and symbol domains, the ambiguity function of OFDM signals with the information-bearing symbols removed exhibits a thumbtack shape, endowing OFDM waveforms with high delay and Doppler resolution, especially when the channel bandwidth is large and/or the coherent processing interval (CPI) is long.
@@
For the above reasons, OFDM is expected to be a competitive technology for ISAC systems. Besides, its compatibility with legacy communication systems makes it a direct and standard-compliant evolutionary solution. For OFDM waveforms used in existing mobile communication systems, numerous studies have demonstrated that sensing can be achieved by using either pilot reference signals [35, 36] or the entire data frame [37] of OFDM signals, highlighting its significant potential for future 6G ISAC scenarios. However, the sensing capabilities of OFDM require further exploration. Some research has focused on optimizing MIMO-OFDM ISAC transmitter designs, such as waveforms and signal optimization [5, 38], interference management [39], power and time-frequency resource allocation [40], and beamforming design [41], but further investigation into efficient receiver signal processing techniques is still required.
@@
To this end, three key challenges must be addressed. First, a unified mathematical model for MIMO-OFDM ISAC is required to reveal the relationship between the system configuration and sensing performance. For instance, increasing the number of subcarriers and symbols can enhance the delay and Doppler resolution of OFDM waveforms, but this also significantly increases the complexity of delay and Doppler estimation. Moreover, joint estimation of target parameters across the spatial, delay, and Doppler domains, significantly escalates signal processing complexity at the receiver, hindering real-time ISAC implementation in practice. Finally, the uniform spherical wave (USW) characteristics in near-field scenarios invalidate traditional far-field algorithms designed based on the conventional uniform plane wave (UPW) model, posing new challenges in computational complexity and estimation performance for extending far-field methods to near-field scenarios. Therefore, a comprehensive and in-depth study of mathematical models and signal processing techniques for MIMO-OFDM ISAC systems is of paramount importance.
@@
Existing relevant overview papers [4, 6, 7, 42] have primarily focused on the fundamental limits, waveform designs, application scenarios, and challenges of generic ISAC, without focusing on MIMO-OFDM ISAC. In [5], the authors considered ISAC signal design and optimization based on OFDM, but signal processing techniques were designed only for far-field scenarios, and joint processing of three-dimensional signals under MIMO-OFDM ISAC was not considered. Unlike prior work, this paper aims to provide a comprehensive and in-depth tutorial overview on MIMO-OFDM ISAC that focuses on signal processing techniques at the receiver side, extending from far-field to near-field scenarios. The main contributions of this work are summarized as follows.
@@
  • First, based on the six basic ISAC modes discussed by 3GPP, we present a unified mathematical ISAC model and characterize the signal transceiving relationships among different ISAC transmitters and receivers. Based on this unified model, various input-output relationships and corresponding signal processing approaches for MIMO-OFDM ISAC systems are elaborated. Moreover, we propose four frameworks for estimating parameters across the spatial, delay, and Doppler domains, including parallel one-domain, sequential one-domain, joint two-domain and joint three-domain parameter estimation. Then, a comprehensive comparison of the four frameworks is provided.
@@
  • Next, we elaborate the four parameter estimation frameworks and corresponding estimation algorithms for far-field scenarios. To this end, the application and performance of various classical radar algorithms are discussed in the context of MIMO-OFDM ISAC, including inverse discrete Fourier transform/discrete Fourier transform (IDFT/DFT)-based methods, subspace-based super-resolution methods, and compressed sensing (CS) methods. Theoretical analysis and numerical results are presented to compare the computational complexity, estimation accuracy, and resolution of these algorithms.
@@
  • Finally, we present target sensing algorithms for the more general near-field scenarios. In particular, we point out the challenges that arise when directly applying far-field based algorithms to near-field sensing problems. Subsequently, we extend the far-field algorithms to near-field sensing and study representative near-field algorithms with low complexity and super resolution. Performance comparisons and analysis of these near-field algorithms are also conducted.
@@
The rest of this paper is organized as follows. Section II introduces the Key Performance Indicators (KPIs) and key technologies of ISAC, as well as the reasons for adopting OFDM waveforms. Section III presents a unified ISAC framework and the corresponding MIMO-OFDM ISAC signal model. Section IV and Section V detail parameter estimation methods for far-field and near-field target sensing, respectively. Section VI provides directions worthy of further investigation, and Section VII concludes this paper.
@@
Notations: Bold face lower- and upper-case letters denote vectors and matrices, respectively. The notations ${\left(\cdot\right)}^{T}$, ${\left(\cdot\right)}^{*}$, ${\left(\cdot\right)}^{H}$ denote transpose, conjugate, and conjugate transpose operations, respectively. The modulus of a complex scalar is given by $|a|$, $|\mathbf{a}|$ and $|\mathbf{A}|$ denote the element-wise modulus of vector $\mathbf{a}$ and matrix $\mathbf{A}$. The transformation of vector $\mathbf{a}$ into a diagonal matrix is represented as $\mathrm{diag}\{\mathbf{a}\}$, and $[\mathbf{a}]_{m}$ denotes the $m$th element of vector $\mathbf{a}$. The operator $\mathrm{vec}\left(\mathbf{A}\right)$ converts matrix $\mathbf{A}$ into a column vector by sequentially stacking its columns. The average of all elements of $\mathbf{A}$ is obtained by $\mathrm{mean}\{\mathbf{A}\}$, and $\mathrm{angle}(a)$ represents the phase of a complex number $a$. The operators ${{\bf{A}}^{\dagger}}$ and $\rm{det}(\mathbf{A})$ denote the pseudoinverse and determinant of matrix $\mathbf{A}$, respectively. The symbols $\otimes$ and $\odot$ denote the Kronecker product and Hadamard product, respectively. The rectangle pulse function is denoted as $\mathrm{rect}(t)=1,t\in\left[0,1\right]$.

### v2

@@ Abstract
Integrated sensing and communication (ISAC) is one of the key usage scenarios for future sixth-generation (6G) mobile communication networks, where communication and sensing (C&S) services are simultaneously provided through shared wireless spectrum, signal processing modules, hardware, and network infrastructure. Such an integration is strengthened by the technology trends in 6G, such as denser network nodes, larger antenna arrays, wider bandwidths, higher frequency bands, and more efficient utilization of spectrum and hardware resources, which incentivize and empower enhanced sensing capabilities. Moreover, emerging applications such as Internet-of-Everything (IoE), autonomous ground and aerial vehicles, virtual reality/augmented reality (VR/AR), and connected intelligence have intensified the demands for both high-quality C&S services, accelerating the development and implementation of ISAC in wireless networks. As in contemporary communication systems, orthogonal frequency-division multiplexing (OFDM) is expected to be the dominant waveform for ISAC, motivating the need for study of both the potential benefits and challenges of OFDM ISAC. Thus, this paper aims to provide a comprehensive tutorial overview of ISAC systems enabled by large-scale multi-input multi-output (MIMO) and OFDM technologies and discuss their fundamental principles, advantages, and enabling signal processing methods. To this end, a unified MIMO-OFDM ISAC system model is first introduced, followed by four frameworks for estimating parameters across the spatial, delay, and Doppler domains, including parallel one-domain, sequential one-domain, joint two-domain, and joint three-domain parameter estimation. Next, sensing algorithms and performance analysis are presented in detail for far-field scenarios where uniform plane wave (UPW) propagation is valid, followed by extensions to near-field scenarios where uniform spherical wave (USW) characteristics must be considered. Finally, the paper presents open challenges and outlines promising avenues for future research on MIMO-OFDM ISAC.
@@ I Introduction
ISAC originated from radar-communication coexistence (RCC) studies [13], which initially focused on interference mitigation for collocated radar and communication systems. With the evolution of wireless communication and digital radar technologies, such as multi-input multi-output (MIMO) [14, 15] and orthogonal frequency division multiplexing (OFDM) [16], the gap between C&S has been gradually bridged. This technological convergence drove the emergence of dual functional radar communication (DFRC) systems [17, 18] that achieve higher spectral efficiency (SE) and hardware integration. To enable the coexistence of radar and communication systems within the same frequency band, a portion of the sub-6 GHz bands was released for radar and communication [19], which promoted the evolution of DFRC towards ISAC. Afterwards, extensive research efforts have been devoted to optimizing ISAC performance by virtue of information theory [20, 21, 22], waveform design [23, 24, 25, 26, 27, 28, 29], signal processing techniques [30, 31], experiments and prototyping [32, 33].
@@
There are three typical design paradigms for ISAC, namely, communication-centric design, sensing-centric design, and C&S co-design. In this paper, we focus on communication-centric designs that exploit communication waveforms to achieve sensing functions. Leveraging globally and densely deployed wireless infrastructure and networks, communication-centric ISAC has emerged as one of the most promising approaches for 6G systems, providing significant advantages over separate C&S designs, as follows.
@@
  • Efficient spectrum utilization: The massive number of Internet-of-Things (IoT) devices and soaring demands for high-speed data services have imposed an unprecedented strain on the limited spectrum resources available for C&S [34]. In conventional DFRC systems, C&S are performed in non-overlapping frequency bands, resulting in fragmented and poor utilization of spectrum resources. Communication-centric ISAC provides an efficient means to improve spectrum efficiency, since the spectrum pool can be flexibly allocated and even completely reused by C&S tasks, as will be discussed in Section II-B.
@@
  • Pervasive sensing via wireless communication networks: Pervasive sensing is an important ISAC vision for 6G that cannot be achieved by traditional radar sensing networks with limited coverage. This issue can be resolved by leveraging existing wireless infrastructure, including cellular base stations (BS) and backhaul links that are densely deployed. Working cooperatively, a collection of wireless networks can form a powerful and pervasive sensor network with ubiquitous coverage and connectivity.
@@
  • Enhanced cost/energy efficiency: Enhancing energy efficiency (EE) and cost efficiency (CE) is an important goal for 6G and ISAC. Pervasive sensing that relies on conventional standalone radars would require the deployment of a large amount of radar equipment, thus degrading EE/CE. Instead, exploiting existing communication systems to perform wireless sensing can significantly reduce the hardware cost and energy consumption for sensing, bringing great value to 6G.
@@
  • High SE: OFDM systems employ inverse fast Fourier transform (IFFT)/ fast Fourier transform (FFT) operations to achieve frequency domain transformations, dividing the channel bandwidth into multiple orthogonal subcarriers for parallel data transmission, thus enabling high SE.
@@
  • Robustness against Inter-Symbol Interference (ISI): The use of a cyclic prefix (CP) in OFDM efficiently mitigates the impact of multipath-induced ISI.
@@
  • Easy equalization: Since OFDM equalization is performed in the frequency domain, the receiver only needs to perform one-tap equalization on each subcarrier, making it much simpler and more efficient than time-domain equalization in single-carrier systems.
@@
  • Flexible time-frequency resource allocation: The number and spacing of subcarriers, together with the power and bit loading on each subcarrier can be dynamically adjusted based on available channel state information (CSI), thereby maximizing network throughput and minimizing error rates while ensuring desired performance for individual users.
@@
For target sensing, OFDM has also been considered to be a promising waveform with the following advantages.
@@
  • Array-manifold-like structure in the subcarrier and symbol domains: After removing random information-bearing symbols, the remaining OFDM signal exhibits an array-manifold-like structure in both the subcarrier and symbol domains, enabling efficient estimation of the delay and Doppler frequency for the sensing targets, as will be shown in detail in Section IV-B.
@@
  • Efficient decoupled estimation of delay and Doppler: The phase variations in OFDM channels induced by delay and Doppler in the subcarrier and symbol domains are independent. This allows for efficient decoupled estimation of delay and Doppler, where the data in the symbol domain can be used as snapshots for delay estimation and vice versa.
@@
  • “Thumbtack-shaped” ambiguity function: In both the subcarrier and symbol domains, the ambiguity function of OFDM signals with the information-bearing symbols removed exhibits a thumbtack shape, endowing OFDM waveforms with high delay and Doppler resolution, especially when the channel bandwidth is large and/or the coherent processing interval (CPI) is long.
Figure 1: Organization of this paper.
@@
For the reasons listed above, and its compatibility with legacy communication systems, OFDM is expected to be the dominant waveform for ISAC. Numerous studies have demonstrated that sensing can be achieved by existing wireless standards using either OFDM pilot reference signals [35, 36] or the entire OFDM symbol [37], providing a hint of its potential for future 6G ISAC scenarios. However, the sensing capabilities of OFDM require further exploration. Some research has focused on optimizing MIMO-OFDM ISAC transmitter designs, such as waveforms [5, 38], interference management [39], power and time-frequency resource allocation [40], and beamforming [41], but further investigation into efficient receiver signal processing techniques is still required.
@@
To this end, three key challenges must be addressed. First, a unified mathematical model for MIMO-OFDM ISAC is required to reveal the relationship between the system configuration and sensing performance. For instance, increasing the number of subcarriers and symbols can enhance the delay and Doppler resolution of OFDM waveforms, but this also increases the complexity of delay and Doppler estimation. Moreover, joint estimation of target parameters across the spatial, delay, and Doppler domains, significantly escalates signal processing complexity at the receiver, hindering real-time ISAC implementation. Finally, the uniform spherical wave (USW) characteristics in near-field scenarios invalidate traditional far-field algorithms designed based on the conventional uniform plane wave (UPW) model, posing new challenges in computational complexity and estimation performance for extending far-field methods to near-field scenarios. Therefore, a comprehensive study of mathematical models and signal processing techniques for MIMO-OFDM ISAC systems is of paramount importance.
@@
Existing relevant overview papers [4, 6, 7, 42] have primarily focused on the fundamental limits, waveform designs, application scenarios, and challenges of generic ISAC, without focusing on MIMO-OFDM designs. In [5], the authors considered ISAC signal design and optimization based on OFDM, but signal processing techniques only for far-field scenarios were considered, and joint processing across the time, frequency, and spatial domains was not considered. Unlike prior work, this paper aims to provide an in-depth tutorial overview of MIMO-OFDM ISAC that focuses on signal processing techniques at the receiver, extending from far-field to near-field scenarios. The main contributions of this work are summarized as follows.
@@
  • First, based on the six basic ISAC modes discussed by 3GPP, we present a unified mathematical ISAC model and characterize the relationships among different ISAC transmitter and receiver designs. Based on this unified model, various input-output relationships and corresponding signal processing approaches for MIMO-OFDM ISAC systems are elaborated. Moreover, we propose four frameworks for estimating parameters across the spatial, delay, and Doppler domains, including parallel one-domain, sequential one-domain, joint two-domain and joint three-domain parameter estimation. Then, a comprehensive comparison of the four frameworks is provided.
@@
  • Next, we elaborate the four parameter estimation frameworks and corresponding estimation algorithms for far-field scenarios. To this end, the application and performance of various classical radar algorithms are discussed in the context of MIMO-OFDM ISAC, including inverse discrete Fourier transform/discrete Fourier transform (IDFT/DFT)-based methods, subspace-based super-resolution methods, and compressed sensing (CS) methods. Theoretical analyses and numerical results are presented to compare the computational complexity, estimation accuracy, and resolution of these algorithms.
@@
  • Finally, we present target sensing algorithms for more general near-field scenarios. In particular, we point out the challenges that arise when directly applying far-field based algorithms to near-field sensing problems. Subsequently, we extend the far-field algorithms to near-field sensing and study representative near-field algorithms with low complexity and super resolution. Performance comparisons and analyses of these algorithms are also conducted.
TABLE I: List of main notations.
  Notations Meanings
  $M_{r}$
Number of antennas at the sensing receiver
$M_{t}$
Number of antennas at the ISAC transmitter
NN Number of OFDM subcarriers
$P$ Number of OFDM symbols
$K$ Number of sensing targets
$\mathbf{q}_{k}(r_{k},\theta_{k})$ Location of the $k$-th target
$r_{k}$ Distance between the $k$-th target and Sen-RX
$\theta_{k}$ Angle from the $k$-th target to Sen-RX
$\tau_{k}$ Delay of the $k$-th target
$\upsilon_{k}$ Doppler frequency shift of the $k$-th target
$\mathbf{x}(t)\in\mathbb{C}^{M_{t}\times 1}$ Signal transmitted by ISAC-TX
$\mathbf{y}_{s}(t)\in\mathbb{C}^{M_{r}\times 1}$ Signal received by Sen-RX
$\mathbf{Y}_{s}\in\mathbb{C}^{M_{r}\times N\times P}$ Sensing signal after FFT demodulation
$\overline{\mathbf{Y}}_{s}\in\mathbb{C}^{M_{r}\times N\times P}$
Sensing signal after element-wise division
$\mathbf{a}_{\rm R}(\theta)\in\mathbb{C}^{M_{r}\times 1}$
Far-field steering vector in the spatial domain
$\mathbf{a}_{\rm R}(r,\theta)\in\mathbb{C}^{M_{r}\times 1}$
Near-field steering vector in the spatial domain
$\mathbf{a}_{\tau}(\tau)\in\mathbb{C}^{N\times 1}$
Steering vector in the delay domain
$\mathbf{a}_{\upsilon}(\upsilon)\in\mathbb{C}^{P\times 1}$
Steering vector in the Doppler domain
$\mathbf{A}_{\theta}\in\mathbb{C}^{M_{r}\times K}$
Far-field steering matrix in the spatial domain
$\mathbf{A}_{\rm R}(r,\theta)\in\mathbb{C}^{M_{r}\times K}$
Near-field steering matrix in the spatial domain
$\mathbf{A}_{\tau}\in\mathbb{C}^{N\times K}$
Steering matrix in the delay domain
$\mathbf{A}_{\upsilon}\in\mathbb{C}^{P\times K}$
Steering matrix in the Doppler domain
$\mathbf{X}_{\theta}\in\mathbb{C}^{M_{r}\times NP}$ Data matrix for AoA estimation
$\mathbf{X}_{\tau}\in\mathbb{C}^{N\times M_{r}P}$ Data matrix for delay estimation
$\mathbf{X}_{\upsilon}\in\mathbb{C}^{P\times M_{r}N}$ Data matrix for Doppler estimation
TABLE II: List of key acronyms.
  Acronyms Definitions
  2D Two-dimensional
3D Three-dimensional
3GPP 3rd Generation Partnership Project
6G Sixth-generation
AoA Angle of arrival
AoD Angle of departure
AWGN Additive white Gaussian noise
BS Base station
C&S Communication and sensing
CS Compressed sensing
CP Cyclic prefix
CPI Coherent processing interval
Com-RX Communication receiver
CKM Channel knowledge map
DFT Discrete Fourier transform
DFRC Dual functional radar communication
ESPRIT
Estimation of signal parameters via
rotational invariance technique
FFT Fast Fourier transform
ISAC Integrated sensing and communication
ISAC-TX ISAC transmitter
ISI Inter-symbol interference
ICI Inter-carrier interference
IDFT Inverse discrete Fourier transform
IFFT Inverse fast Fourier transform
LoS Line-of-sight
LS Least squares
MIMO Multi-input multi-output
MUSIC Multiple signal classification
MRC Maximum-ratio combining
MMSE Minimum mean square error
NLoS Non-line-of-sight
NOMA Non-orthogonal multiple access
OFDM Orthogonal frequency division multiplexing
OTFS Orthogonal time frequency space
OMP Orthogonal matching pursuit
PAPR Peak-to-average power ratio
PM Propagator method
RMSE Root mean square error
SI Self-interference
SE Spectral efficiency
Sen-RX Sensing receiver
SNR Signal-to-noise ratio
TLS Total least square
UPW Uniform plane wave
USW Uniform spherical wave
ULA Uniform linear array
UE User equipment
XL-MIMO Extremely large-scale multi-input multi-output
ZF Zero-forcing
@@
As outlined in Fig. 1, the rest of the paper is organized as follows. Section II introduces the Key Performance Indicators (KPIs) and key technologies of ISAC, as well as the reasons for adopting OFDM waveforms. Section III presents a unified ISAC framework and the corresponding MIMO-OFDM ISAC signal model. Section IV and Section V detail parameter estimation methods for far-field and near-field target sensing, respectively. Section VI provides directions worthy of further investigation, and Section VII concludes this paper.
@@
Notation: Bold face lower- and upper-case letters denote vectors and matrices, respectively. The notations ${\left(\cdot\right)}^{T}$, $\left(\cdot\right)$$^{*}$, ${\left(\cdot\right)}^{H}$ denote transpose, conjugate, and conjugate transpose operations, respectively. The modulus of a complex scalar is given by $|a|$, $|\mathbf{a}|$ and $|\mathbf{A}|$ denote the element-wise modulus of vector $\mathbf{a}$ and matrix $\mathbf{A}$. The transformation of vector $\mathbf{a}$ into a diagonal matrix is represented as $\mathrm{diag}\{\mathbf{a}\}$, and $\mathbf{a}]_{m}$ denotes the $m$th element of vector $\mathbf{a}$. The operator $\mathrm{vec}\left(\mathbf{A}\right)$ converts matrix $\mathbf{A}$ into a column vector by sequentially stacking its columns. The average of all elements of $\mathbf{A}$ is obtained by $\mathrm{mean}\{\mathbf{A}\}$, and $\mathrm{angle}(a)$ represents the phase of a complex number $a$. The operators $\bf{A}$$^{\dagger}$} and $\rm{det}(\mathbf{A})$ denote the pseudoinverse and determinant of matrix $\mathbf{A}$, respectively. The symbols $\otimes$ and $\odot$ denote the Kronecker product and Hadamard product, respectively. The rectangle pulse function is denoted as $\mathrm{rect}(t)=1,t\in\left[0,1\right]$. Notations frequently used in this paper are given in Table I, and Table II summarizes key acronyms.